# Leveraging the Power of MLLMs for Gloss-Free Sign Language Translation

### v1

@@ Abstract
Sign language translation (SLT) is a challenging task that involves translating sign language images into spoken language. For SLT models to perform this task successfully, they must bridge the modality gap and identify subtle variations in sign language components to understand their meanings accurately. To address these challenges, we propose a novel gloss-free SLT framework called Multimodal Sign Language Translation (MMSLT), which leverages the representational capabilities of off-the-shelf multimodal large language models (MLLMs). Specifically, we generate detailed textual descriptions of sign language components using MLLMs. Then, through our proposed multimodal-language pre-training module, we integrate these description features with sign video features to align them within the spoken sentence space. Our approach achieves state-of-the-art performance on benchmark datasets PHOENIX14T and CSL-Daily, highlighting the potential of MLLMs to be effectively utilized in SLT.
**footnotetext: Equally contributed\$$\dagger$\$$\dagger$\$footnotetext: Correspoinding author
@@ 1 Introduction
Sign language is the primary mode of communication for deaf individuals. It relies on visual elements such as hand gestures, body movements, and facial expressions [47, 46, 44], collectively called SL components. The sign language translation (SLT) task aims to bridge communication gaps by converting sign language videos into spoken sentences. However, this task poses significant challenges, as the model needs to identify SL components and convert them into spoken sentences. Specifically, it requires overcoming the modality gap—transitioning from visual cues to text—while also understanding the intricate cross-modal relationships involved [52, 16, 66].
Figure 1: Comparison of the proposed MMSLT with GFSLT-VLP [67], a representative gloss-free SLT approach. Previous gloss-free SLT methods extract visual features from sign language images and align them with spoken sentences. In contrast, our MMSLT leverages an MLLM to represent sign language images as text, integrates the resulting representations with the images, and aligns the fused modalities with spoken sentences.
To enhance SLT, gloss-based models [5, 53, 63] that use mid-level supervision with glosses have been developed, as well as weakly supervised gloss-free models [62, 66, 27] that indirectly leverage glosses to extract visual features through an encoder trained on sign language recognition datasets. However, these approaches still rely on labor-intensive gloss annotations, which pose scalability challenges and create information bottlenecks [23, 24]. Therefore, recent studies have shifted towards gloss-free SLT models that directly translate sign language videos into spoken language [67, 57, 18]. Most gloss-free models utilize modules such as visual encoders and text decoders to extract visual features from sign language images, transforming these features into text-like representations that large language models (LLMs) can process for translation [67, 1, 18], as illustrated in Fig. 1. However, since these visual features are obtained from sign language images, they can be affected by visual information unrelated to SLT, such as background elements or clothing colors, limiting their ability to accurately represent SL components.
@@
Inspired by these observations, we propose a novel gloss-free approach called Multimodal Sign Language Translation framework (MMSLT). Our MMSLT consists of two main modules to enable a comprehensive understanding of sign language through the complementary integration of modalities, SL descriptions and sign language images, as shown in Fig. 4. First, the Generating Sign Language Description via MLLM (GSD-MLLM) module prompts a pre-trained MLLM to generate SL descriptions. Then, the Multimodal-Language Pre-training (MMLP) module integrates the two modality features and aligns them with the target spoken sentence space, bridging the modality gap and improving translation accuracy. Furthermore, to enhance efficiency, we introduce a description mapper that predicts description embeddings from the embedding features of sign language images, alleviating the computational burden of using MLLM during each inference. To evaluate MMSLT, we perform extensive experiments on two benchmark datasets: PHOENIX14T [5] and CSL-Daily [68]. The results demonstrate that MMSLT significantly outperforms previous state-of-the-art (SOTA) gloss-free SLT methods, particularly on the large-scale CSL-Daily dataset. The main contributions of MMSLT are as follows:
@@
  • By analyzing various MLLMs and prompts, we propose the GSD-MLLM module, which generates detailed SL descriptions from sign language images.
  • We introduce the MMLP module, which effectively integrates two modalities, SL descriptions and sign language images, and aligns them with the target sentence space to reduce the modality gap.
@@ 2.1 Gloss-free Sign Language Translation
Glosses provide intermediate supervision by expressing SL components in textual form, which is crucial for understanding sign language and translating it into spoken language [6, 53]. In contrast, gloss-free SLT bypasses glosses and directly translates sign language videos into spoken sentences using only paired video-sentence data. Consequently, this approach tends to have a greater modality gap than gloss-based [6, 53, 69, 68, 61, 28, 8, 9, 63] or weakly supervised gloss-free SLT models [30, 62, 17, 27], generally resulting in lower performance. To address this gap, CSGCR [66] predicts possible language words, generates sentences based on these predictions, and selects the most appropriate sentence using cosine similarity with sign language images. Recently, GFSLT-VLP [67] introduced a pre-training method that aligns sign language images with spoken sentences, converting them into text-like features familiar to LLMs.
Leveraging this advancement, subsequent gloss-free models focus on generating more comprehensible features for LLMs. FLa-LLM [12] introduces a two-step approach: initially training on visual information from sign language images using a lightweight translation model and fine-tuning the LLM for SLT. Sign2GPT [57] pre-trains a sign encoder by aligning visual features with prototypes under the supervision of pseudo-glosses—words from spoken sentences via Parts-of-Speech (POS) tagging (e.g., nouns and numbers)—and then utilizes it for SLT. Lastly, SignLLM [18] normalizes input video to generate sign tokens with linguistic features compatible with LLMs. However, to date, no SLT study has directly represented sign language images in a textual format. Representing the same data across different modalities, as proposed in our approach, allows the two modalities to complement each other and create synergy. Furthermore, this approach can also help mitigate the modality gap.
@@ 2.3 Vision-Language Pre-training
GFSLT-VLP [67] is the first study to utilize CLIP [43], a pioneering approach in VLP, to tackle the modality gap in gloss-free SLT. Like CLIP, GFSLT-VLP adopts a dual-stream structure, pre-training the alignment between sign videos and spoken sentences before performing the translation. However, this approach has limitations in providing frame-level supervision. Therefore, VAP [27] proposes a method to learn the alignment between frames in sign language videos and words in spoken sentences. Despite these advancements, existing SLT studies using VLP have primarily focused on elucidating the relationship between a single modality (visual) and a single modality (text). As far as we know, no SLT research currently utilizes multimodality (visual and text) as input to learn the alignment for a single modality (text).
@@ 3 MLLMs and Prompts: Preliminary Analysis
Video-based MLLM vs Image-based MLLM. We first examine the potential of video-based and image-based MLLMs for SLT by evaluating their ability to describe SL components. As shown in Fig. 2(a), video-based MLLMs, such as Video-LLaMA [65] and Video-LLaVA [34], demonstrate limited capability in capturing SL components. Instead, they provide simplistic summaries, like ‘using her hands to gesture’ or repeat the same content, and include general descriptions of situations unrelated to sign language, such as ‘wearing a suit’ or inaccurate information. On the other hand, image-based MLLMs such as LLaVa-Next [31], InternVL [11], QwenVL2 [55], Pixtral [2], and LLaVA-OneVision [29] generate detailed descriptions of SL components, such as ‘fingers slightly spread apart,’ or ‘eyes are focused and directed toward.’ Notably, as shown in Fig. 2(c), LLaVA-OneVision provides descriptions focused solely on SL components without including unrelated information. In this experiment, large models, such as LLaVa-OneVision 72B, are excluded from the analysis due to the processing time and resource constraints associated with generating SL descriptions.
Prompt Engineering. To effectively leverage MLLMs, it is crucial to select an appropriate model and formulate prompts that explicitly delineate the information to be extracted or provide contextual clarity for the content. To achieve this, we perform inference with six distinct prompts, as shown in Fig. 3. These can be grouped into simple (1, 2), detailed (3, 4), and in-context (5, 6) types. Prompts (1), (2), and (4) elicit responses that primarily focus on irrelevant information, as shown in Fig. 2(b), while prompts (3), (5), and (6) generate responses more concentrated on SL components. However, responses to prompt (6) contain inaccuracies, while the responses to prompt (5) are more accurate and provide more detailed information. Nevertheless, they often reuse example sentences and produce repetitive answers across consecutive frames. In contrast, as shown in Fig. 2(c), prompt (3) provides detailed descriptions of SL components, such as hand shape, signer’s gaze, and mouth shape, effectively conveying the meanings of the signer’s gestures and facial expressions. Furthermore, it avoids referencing external information irrelevant to understanding sign language and successfully captures subtle differences between consecutive sign gestures.
@@ 4 Method
Figure 4: MMSLT overview. MMSLT comprises two modules: (a) Generating Sign Descriptions via MLLM (GSD-MLLM): This module utilizes prompts along with sign language videos as inputs to the MLLM, generating SL descriptions that describe the SL components (dashed line in the figure). (b) Multimodal Sign Language Pre-training (MMLP): In this module, sign videos and SL descriptions are integrated and aligned with spoken sentences. We introduce a description mapper, which allows efficient inference without requiring the MLLM by approximating the embedding features of SL descriptions based on the sign language video (solid line in the figure). Following the MMLP, pre-trained networks are fine-tuned, and the initialized LLM decoder is trained for SLT.
This section introduces our proposed MMSLT framework. We begin with a detailed explanation of how SL descriptions are generated using MLLM in Sec. 4.1. Next, Sec. 4.2 describes the MMLP module, which integrates sign language images and the corresponding SL descriptions while aligning them with the target spoken sentences. Finally, Sec. 4.3 explains the training process of SLT.
@@ 4.1 Generating SL Descriptions via MLLM
Based on the analysis in Sec. 3, we utilize an image-based MLLM with a prompt $p$ to generate a set of SL descriptions $d_{i}=\{d_{i,t}\}_{t=1}^{T}$ for the $i$-th sign language video ${SV}_{i}\in\mathbb{R}^{T\times H\times W}$, as described in Fig. 4. In this context, the total number of sign language videos is $N$, the height and width of the frames are $H$ and $W$, respectively. The total number of frames is $T$. However, we observe that the MLLM has limitations in accurately capturing the facial expressions of signers, often producing fragmented descriptions of lip shapes, such as “smiling,” or incorrectly depicting closed eyes as “open eyes.” This reveals that although SL descriptions capture SL components, they are sometimes inaccurate and lack detail.
@@ 4.2 Multimodal-Language Pre-training
To facilitate synergy between two modalities, we integrate sign language images into the SL descriptions. It is evident that a modality gap exists between sign language videos and spoken sentences. To bridge this gap, we propose MMLP, a pre-training module that effectively fuses sign language videos with their corresponding SL descriptions and learns to align them with spoken sentences, as illustrated in Fig. 4.
Description Mapper. First, we extract the visual feature $V_{j}\in\mathbb{R}^{T\times C}$ from the $j$-th sign language video ${SV}_{j}$ within a mini-batch of size $B$ using the image encoder $\psi_{vis}$, which is a ResNet18 [20] pre-trained on ImageNet [14], where $C$ represents the dimensionality of $V_{j}$. Simultaneously, we obtain the description embedding feature $D_{j,t}\in\mathbb{R}^{1\times\bar{C}}$ from the SL description of each frame $d_{j,t}$ using the description encoder $\psi_{de}$, a pre-trained, frozen 12-layer BERT [15]. In this context, we define the embedding feature of the [CLS] token as $D_{j,t}$, encoding sentence-level information, with $\bar{C}$ denoting the dimensionality of $D_{j,t}$. These description embedding features collectively form the description feature $D_{j}=\{D_{j,t}\}_{t=1}^{T}$. However, if $D_{j}$ is directly utilized to integrate sign language videos and SL descriptions, generating SL descriptions via MLLM becomes necessary during the inference. This necessity leads to increased computational costs and extended inference times. To address this issue, we propose a description mapper $\psi_{dm}$ characterized by a simple two-layer MLP structure that predicts $D_{j}$ from $V_{j}$. This also helps bridge the modality gap between the input data sources and allows us to obtain the approximated description feature $\hat{D}_{j}$, which can be expressed as follows:
@@
Modality Adapter. To combine the visual feature and the approximated SL description feature, we propose a modality adapter $\psi_{ma}$ consisting of a 1D convolution, a max-pooling layer, and a two-layer MLP. Since sign language typically consists of multiple frames, and $V_{j}$ and $\hat{D}_{j}$ are derived from consecutive sign language frames, we employ a 1D convolution for temporal modeling consistent with previous studies [67, 12, 57]. Additionally, MLP layers are employed to integrate information from both modalities. The captured $V_{j}$ and $\hat{D}_{j}$ are concatenated and then input into the modality adapter, resulting in the extraction of the sign element feature ${SE}_{j}\in\mathbb{R}^{T^{\prime}\times C^{\prime}}$, where $T^{\prime}$ is the reduced sequence length, and $C^{\prime}$ is the embedding dimension. This process can be expressed as follows:
@@
Multimodal Encoder with LoRA. To extract the representation of ${SE}_{i}$, we introduce a multimodal encoder $\psi_{enc}$. For this purpose, we utilize the mBART encoder [48], which consists of 12 layers and is initialized with parameters pre-trained on a large corpus. To facilitate adaptation to the sign language dataset while preserving the pre-trained knowledge, we apply the LoRA [22] technique, which enables parameter-efficient fine-tuning by adding low-rank matrices into the weight matrix of the LLM. The resulting output, referred to as the multimodal visual-textual feature (multimodal feature), is denoted as ${M}_{j}\in\mathbb{R}^{T^{\prime}\times C^{\prime}}$ and can be represented as follows:
@@
Multimodal-Language Alignment. ${M}_{j}$ encompasses visual and textual information; however, a modality gap still remains with the target spoken sentence. To tackle this challenge and align ${M}_{j}$ with the target spoken sentence, we propose a multimodal-language alignment method.
Let ${SL}_{j}$ be the target spoken sentence corresponding to ${SV}_{j}$. We embed ${SL}_{j}$ into text features $L_{j}\in\mathbb{R}^{\bar{T}\times C^{\prime}}$, where $\bar{T}$ denotes the number of tokens in the spoken sentence, using a text encoder $\psi_{te}$, a frozen 12-layer mBART encoder [48] pre-trained on a large corpus, similar to the $\psi_{enc}$. Since our objective is to align $\{M_{j},L_{j}\}_{j=1}^{B}$, we apply contrastive learning, which is consistent with previous studies [67, 27]. Specifically, we perform average pooling over the frame sequence $T^{\prime}$ and the token sequence $\bar{T}$ to obtain the global multimodal feature $\tilde{M}_{j}\in\mathbb{R}^{C^{\prime}}$ and global spoken-sentence feature $\tilde{L}_{j}\in\mathbb{R}^{C^{\prime}}$. Finally, we align the pairs $\{\tilde{M}_{j},\tilde{L}_{j}\}_{j=1}^{N}$ using the loss function as follows:

### v2

@@ Abstract
Sign language translation (SLT) is a challenging task that involves translating sign language images into spoken language. For SLT models to perform this task successfully, they must bridge the modality gap and identify subtle variations in sign language components to understand their meanings accurately. To address these challenges, we propose a novel gloss-free SLT framework called Multimodal Sign Language Translation (MMSLT), which leverages the representational capabilities of off-the-shelf multimodal large language models (MLLMs). Specifically, we use MLLMs to generate detailed textual descriptions of sign language components. Then, through our proposed multimodal-language pre-training module, we integrate these description features with sign video features to align them within the spoken sentence space. Our approach achieves state-of-the-art performance on benchmark datasets PHOENIX14T and CSL-Daily, highlighting the potential of MLLMs to be utilized effectively in SLT. Code is available at https://github.com/hwjeon98/MMSLT.
footnotetext: *These authors contributed equally.footnotetext: Corresponding author.
@@ 1 Introduction
Sign language (SL) is the primary mode of communication for deaf individuals. It relies on visual elements such as hand gestures, body movements, and facial expressions [47, 46, 44], collectively called SL components. The sign language translation (SLT) task aims to bridge communication gaps by converting SL videos into spoken sentences. However, this task poses significant challenges, as the model needs to identify SL components and convert them into spoken sentences. Specifically, it requires overcoming the modality gap—transitioning from visual cues to text—while also understanding the intricate cross-modal relationships involved [52, 16, 66].
Figure 1: Comparison of the proposed MMSLT with GFSLT-VLP [67], a representative gloss-free SLT approach. Previous gloss-free SLT methods extract visual features from SL images and align them with spoken sentences. In contrast, our MMSLT leverages an MLLM to represent SL images as text, integrates the resulting representations with the images, and aligns the fused modalities with spoken sentences.
To enhance SLT, gloss-based models [5, 53, 63] that use mid-level supervision with glosses have been developed, as well as weakly supervised gloss-free models [62, 66, 27] that indirectly leverage glosses to extract visual features through an encoder trained on SL recognition datasets. However, these approaches still rely on labor-intensive gloss annotations, which pose scalability challenges and create information bottlenecks [23, 24]. Therefore, recent studies have shifted towards gloss-free SLT models that directly translate SL videos into spoken language [67, 57, 18]. Most gloss-free models utilize modules such as visual encoders and text decoders to extract visual features from SL images, transforming these features into text-like representations that large language models (LLMs) can process for translation [67, 1, 18], as illustrated in Fig. 1. However, since these visual features are obtained from SL images, they can be affected by visual information unrelated to SLT, such as background elements or clothing colors, limiting their ability to accurately represent SL components.
@@
Inspired by these observations, we propose a novel gloss-free approach called Multimodal Sign Language Translation framework (MMSLT). Our MMSLT consists of two main modules to enable a comprehensive understanding of SL through the complementary integration of modalities, SL descriptions and SL images, as shown in Fig. 4. First, the Generating Sign Language Description via MLLM (GSD-MLLM) module serves as a pre-processing step, prompting a pre-trained MLLM to generate SL descriptions. Then, the Multimodal-Language Pre-training (MMLP) module integrates the two modality features and aligns them with the target spoken sentence space, bridging the modality gap and improving translation accuracy. Furthermore, to enhance efficiency, we introduce a description mapper that predicts description embeddings from the embedding features of SL images, alleviating the computational burden of using MLLM during each inference.
To evaluate MMSLT, we perform extensive experiments on two benchmark datasets: PHOENIX14T [5] and CSL-Daily [68]. Results show that MMSLT significantly outperforms previous state-of-the-art (SOTA) gloss-free SLT methods, particularly on the large-scale CSL-Daily dataset. The main contributions of MMSLT are as follows:
@@
  • By analyzing various MLLMs and prompts, we propose the GSD-MLLM module as a pre-processing step, which generates detailed SL descriptions from SL images.
  • We introduce the MMLP module, which effectively integrates two modalities, SL descriptions and SL images, and aligns them with the target sentence space to reduce the modality gap.
@@ 2.1 Gloss-free Sign Language Translation
Glosses provide intermediate supervision by expressing SL components in textual form, which is crucial for understanding SL and translating it into spoken language [6, 53]. In contrast, gloss-free SLT bypasses glosses and directly translates SL videos into spoken sentences using only paired video-sentence data. Consequently, this approach tends to have a greater modality gap than gloss-based [6, 53, 69, 68, 61, 28, 8, 9, 63] or weakly supervised gloss-free SLT models [30, 62, 17, 27], which generally results in lower performance. To address this gap, CSGCR [66] predicts possible language words, generates sentences based on these predictions, and selects the most appropriate sentence using cosine similarity with SL images. GFSLT-VLP [67] introduces a pre-training method that aligns SL images with spoken sentences, converting them into LLM-friendly features.
Leveraging this advancement, subsequent gloss-free models focus on generating more comprehensible features for LLMs. FLa-LLM [12] introduces a two-step approach: initially training on visual information from SL images using a lightweight translation model and fine-tuning the LLM for SLT. Sign2GPT [57] pre-trains a sign encoder by aligning visual features with prototypes under the supervision of pseudo-glosses—words from spoken sentences via Parts-of-Speech (POS) tagging (e.g., nouns and numbers)—and then utilizes it for SLT. Lastly, SignLLM [18] normalizes input video to generate sign tokens with linguistic features compatible with LLMs. However, to date, no SLT study has directly represented SL images in a textual format. Representing the same data across different modalities, as proposed in our approach, allows the two modalities to complement each other and create synergy. Furthermore, this approach can also help mitigate the modality gap.
@@ 2.3 Vision-Language Pre-training
GFSLT-VLP [67] is the first study to utilize CLIP [43], a pioneering approach in VLP, to tackle the modality gap in gloss-free SLT. Like CLIP, GFSLT-VLP adopts a dual-stream structure, pre-training the alignment between sign videos and spoken sentences before performing the translation. However, this approach has limitations in providing frame-level supervision. Therefore, VAP [27] proposes a method to learn the alignment between frames in SL videos and words in spoken sentences. Despite these advancements, existing SLT studies using VLP have primarily focused on elucidating the relationship between a single modality (visual) and a single modality (text). As far as we know, no SLT research currently utilizes multimodality (visual and text) as input to learn the alignment for a single modality (text).
@@ 3 MLLMs and Prompts: Preliminary Analysis
Video-based MLLM vs Image-based MLLM. We first examine the potential of video-based and image-based MLLMs for SLT by evaluating their ability to describe SL components. As shown in Fig. 2(a), video-based MLLMs, such as Video-LLaMA [65] and Video-LLaVA [34], demonstrate limited capability in capturing SL components. Instead, they provide simplistic summaries, like ‘using her hands to gesture’ or repeat the same content, and include general descriptions of situations unrelated to SL, such as ‘wearing a suit’ or inaccurate information. On the other hand, image-based MLLMs such as LLaVA-Next [31], InternVL [11], QwenVL2 [55], Pixtral [2], and LLaVA-OneVision [29] generate detailed descriptions of SL components, such as ‘fingers slightly spread apart,’ or ‘eyes are focused and directed toward.’ Notably, as shown in Fig. 2(c), LLaVA-OneVision provides descriptions focused solely on SL components without including unrelated information. In this experiment, large models, such as LLaVA-OneVision 72B, are excluded from the analysis due to the processing time and resource constraints associated with generating SL descriptions.
Prompt Engineering. To effectively leverage MLLMs, it is crucial to select an appropriate model and formulate prompts that explicitly delineate the information to be extracted or provide contextual clarity for the content. To achieve this, we perform inference with six distinct prompts, as shown in Fig. 3. These can be grouped into simple (1, 2), detailed (3, 4), and in-context (5, 6) types. Prompts (1), (2), and (4) elicit responses that primarily focus on irrelevant information, as shown in Fig. 2(b), while prompts (3), (5), and (6) generate responses more concentrated on SL components. However, responses to prompt (6) contain inaccuracies, while the responses to prompt (5) are more accurate and provide more detailed information. Nevertheless, they often reuse example sentences and produce repetitive answers across consecutive frames. In contrast, as shown in Fig. 2(c), prompt (3) provides detailed descriptions of SL components, such as hand shape, signer’s gaze, and mouth shape, effectively conveying the meanings of the signer’s gestures and facial expressions. Furthermore, it avoids referencing external information irrelevant to understanding SL and successfully captures subtle differences between consecutive sign gestures.
@@ 4 Method
Figure 4: MMSLT overview. MMSLT comprises two modules: (a) Generating Sign Descriptions via MLLM (GSD-MLLM): This module utilizes prompts along with SL videos as inputs to the MLLM, generating SL descriptions that describe the SL components (dashed line in the figure). (b) Multimodal Sign Language Pre-training (MMLP): In this module, sign videos and SL descriptions are integrated and aligned with spoken sentences. We introduce a description mapper, which allows efficient inference without requiring the MLLM by approximating the embedding features of SL descriptions based on the SL video (solid line in the figure). Following the MMLP, pre-trained networks are fine-tuned, and the initialized LLM decoder is trained for SLT.
This section introduces our proposed MMSLT framework. We begin with a detailed explanation of how SL descriptions are generated using MLLM in Sec. 4.1. Next, Sec. 4.2 describes the MMLP module, which integrates SL images and the corresponding SL descriptions while aligning them with the target spoken sentences. Finally, Sec. 4.3 explains the training process of SLT.
@@ 4.1 Generating SL Descriptions via MLLM
Based on the analysis in Sec. 3, we utilize an image-based MLLM with a prompt $p$ to generate a set of SL descriptions $d_{j}=\{d_{j,t}\}_{t=1}^{T}$ for the $j$-th SL video ${SV}_{j}\in\mathbb{R}^{T\times H\times W}$, as described in Fig. 4. In this context, the total number of SL videos is $N$, the height and width of the frames are $H$ and $W$, respectively. The total number of frames is $T$.
@@ 4.2 Multimodal-Language Pre-training
To facilitate synergy between two modalities, we integrate SL images into the SL descriptions. It is evident that a modality gap exists between SL videos and spoken sentences. To bridge this gap, we propose MMLP, a pre-training module that effectively fuses SL videos with their corresponding SL descriptions and learns to align them with spoken sentences, as illustrated in Fig. 4.
Description Mapper. First, we extract the visual feature $V_{j}\in\mathbb{R}^{T\times C}$ from the $j$-th SL video ${SV}_{j}$ within a mini-batch of size $B$ using the image encoder $\psi_{vis}$, which is a ResNet18 [20] pre-trained on ImageNet [14], where $C$ represents the dimensionality of $V_{j}$. Simultaneously, we obtain the description embedding feature $D_{j,t}\in\mathbb{R}^{1\times\bar{C}}$ from the SL description of each frame $d_{j,t}$ using the description encoder $\psi_{de}$, a pre-trained, frozen 12-layer BERT [15]. In this context, we define the embedding feature of the [CLS] token as $D_{j,t}$, encoding sentence-level information, with $\bar{C}$ denoting the dimensionality of $D_{j,t}$. These description embedding features collectively form the description feature $D_{j}=\{D_{j,t}\}_{t=1}^{T}$. However, if $D_{j}$ is directly utilized to integrate SL videos and SL descriptions, generating SL descriptions via MLLM becomes necessary during the inference. This necessity leads to increased computational costs and extended inference times. To address this issue, we propose a description mapper $\psi_{dm}$ characterized by a simple two-layer MLP structure that predicts $D_{j}$ from $V_{j}$. This also helps bridge the modality gap between the input data sources and allows us to obtain the approximated description feature $\hat{D}_{j}$, which can be expressed as follows:
@@
Modality Adapter. To combine the visual feature and the approximated SL description feature, we introduce a modality adapter $\psi_{ma}$ consisting of two successive 1D convolution and max-pooling layers, followed by a two-layer MLP. Since SL typically consists of multiple frames, and $V_{j}$ and $\hat{D}_{j}$ are derived from consecutive SL frames, we employ a 1D convolution for temporal modeling consistent with previous studies [67, 12, 57]. Additionally, MLP layers are employed to integrate information from both modalities. The captured $V_{j}$ and $\hat{D}_{j}$ are concatenated and then input into the modality adapter, resulting in the extraction of the sign element feature ${SE}_{j}\in\mathbb{R}^{T^{\prime}\times C^{\prime}}$, where $T^{\prime}$ is the reduced sequence length, and $C^{\prime}$ is the embedding dimension. This process can be expressed as follows:
@@
Multimodal Encoder with LoRA. To extract the long-term representation of ${SE}_{i}$ [67], we introduce a multimodal encoder $\psi_{enc}$. For this purpose, we utilize the mBART encoder [48], which consists of 12 layers and is initialized with parameters pre-trained on a large corpus. To facilitate adaptation to the SL dataset while preserving the pre-trained knowledge, we apply the LoRA [22] technique, which enables parameter-efficient fine-tuning by adding low-rank matrices into the weight matrix of the LLM. The resulting output, referred to as the multimodal visual-textual feature (multimodal feature), is denoted as ${M}_{j}\in\mathbb{R}^{T^{\prime}\times C^{\prime}}$ and can be represented as follows:
@@
Multimodal-Language Alignment. ${M}_{j}$ encompasses visual and textual information; however, a modality gap still remains with the target spoken sentence. To tackle this challenge and align ${M}_{j}$ with the target spoken sentence, we present a multimodal-language alignment method.
Let ${SL}_{j}$ be the target spoken sentence corresponding to ${SV}_{j}$. We embed ${SL}_{j}$ into text features $L_{j}\in\mathbb{R}^{\bar{T}\times C^{\prime}}$, where $\bar{T}$ denotes the number of tokens in the spoken sentence, using a text encoder $\psi_{te}$, a frozen 12-layer mBART encoder [48] pre-trained on a large corpus, similar to the $\psi_{enc}$. Since our objective is to align $\{M_{j},L_{j}\}_{j=1}^{B}$, we apply contrastive learning, which is consistent with previous studies [67, 27]. Specifically, we perform average pooling over the frame sequence $T^{\prime}$ and the token sequence $\bar{T}$ to obtain the global multimodal feature $\tilde{M}_{j}\in\mathbb{R}^{C^{\prime}}$ and global spoken-sentence feature $\tilde{L}_{j}\in\mathbb{R}^{C^{\prime}}$. Finally, we align the pairs $\{\tilde{M}_{j},\tilde{L}_{j}\}_{j=1}^{B}$ using the loss function as follows: